# Optimistic Verifiable Training by Controlling Hardware Nondeterminism

### 14 Mar 2024 v1

@@ 5.1 Accumulation Errors Start at Higher Precision Bits
Our key idea is that if nondeterminism of training between GPU types occurs due to floating-point operations, then any error will initially be introduced in the lower bits. Therefore, if both $\mathrm{trainer}$ and $\mathrm{auditor}$ train at a higher floating-point (e.g. $b_{tr}=64$) precision than the client’s target model precision (e.g. $b_{m}=32$) and then periodically round (e.g. $b_{r}=32$) after intermediate computation steps (e.g. a convolution layer, if operating on a per-layer abstraction such as ), they can help “erase” any errors from accumulating due to nondeterminism. Table 1 (Row 1) provides a toy example of how error in higher precision arithmetic is erased when rounding the result to lower precision.
@@
Unfortunately, simply rounding to the nearest FP32 after each computation during training is not sufficient for eliminating nondeterminism due to divergence spanning the rounding boundary. Consider Case A in Figure 4, which shows a divergence in the output of a computation using FP64 on two different GPUs. Because the outputs of GPU 1 and 2 are on different sides of the boundary, rounding to the nearest FP32 results in rounding to different values, introducing error. Table 1 (Row 2) provides a toy example of how error in higher precision arithmetic might transfer to the lower precision rounded result.
@@
What if the $\mathrm{trainer}$ records their rounding choice (e.g. up, down, none) for every intermediate computation? The $\mathrm{auditor}$ could then copy the trainer’s choice, and therefore round to the exact same value and successfully control for nondeterminism. However, as we show in Case B in Figure 4, the $\mathrm{auditor}$ should not copy the $\mathrm{trainer}$’s behavior for every output. If a computation output on GPU 1 is too close to the rounded value, then it is possible that GPU 2 is also close in distance but from the opposite direction – in this case, the $\mathrm{auditor}$ should ignore the $\mathrm{trainer}$’s choice. We therefore need to introduce a threshold $\tau$ under which the $\mathrm{trainer}$ does not record their rounding choice.
@@
The success of this method relies on our ability to upper bound the divergence $d_{div}$ between any two different GPUs for any intermediate computation $f$ (i.e. difference in outputs for the same input). Let $\epsilon_{b}$ represent the distance between two FP32 values, after rounding to $b_{r}$ bits of the mantissa (Figure 3) and controlling for the exponent. We need to select $b_{r}$ and $\tau$ such that $d_{div}<\epsilon_{b_{r}}$ and $d_{div}<2\tau$ (Figure 4). Because the set of possible floating point numbers is finite, there do exist optimal bounds for $b_{r}$ and $\tau$ – in practice, we find that $b_{r}\leq 32$ and $\tau>0.25*\epsilon_{32}$ are sufficient for standard intermediate computations in neural network training (e.g. convolution, layer norm) in FP64, and experiment with different values for $b_{r}$ in Section 6.
@@
We next discuss the specific functions our method requires, algorithm details for both the $\mathrm{trainer}$ and $\mathrm{auditor}$, and our approach for reducing storage cost.
@@ 5.5 Reducing storage cost
Adaptive Threshold: Recall that we need to select a threshold $\tau$ that controls for whether the $\mathrm{trainer}$ logs a rounding choice, or instead logs 1 which the $\mathrm{auditor}$ ignores. The more one can increase $\tau$, the more 1 values are recorded which, if significant, can make the rounding more compressible (due to an increase in long sequences of 1s). Furthermore, it is possible that the divergence $d_{div}$ between outputs on two different GPUs, given the same input, is function-specific. For example, while the convolution operation requires several matrix multiplications that might result in large floating point accumulation error, operations such as scaling are unlikely to result in large $d_{div}$, and a larger $\tau$ can be applied. We develop an efficient algorithm (Algorithm 4 in the Appendix) to find the optimal value for $\tau$ given a particular layer and data of output values that led to different rounding choices between any two GPUs (e.g. Case A in Figure 4). For a given rounding amount $b_{r}$ and training precision $b_{tr}$, the algorithm performs a binary search between $\tau=0.25*\epsilon_{32}$ (our assumed upper bound for the $d_{div}$ between two GPUs for any function) and $\tau=0.5*\epsilon_{b_{r}}$ (the rounding boundary). By performing this procedure for the different intermediate computations in a model, the $\mathrm{trainer}$ can hope to log more values of 1, leading to a stronger compression of the rounding log $F$.
@@ 6 Empirical Results
We evaluate our verifiable training method on the two large-scale models listed below with all possible $\mathrm{trainer}$ and $\mathrm{auditor}$ pairs across the following three NVIDIA GPU architectures: A40, TITAN Xp, and RTX 2080 Ti (see Appendix Section A for more details). Additionally, in Section 6.4, we compare our method with a proof-based system for logistic regression as no known proof-based systems exist for the large-scale models we have considered.
@@
In Figure 2 we show that nondeterminism due to GPU architecture exists for both training tasks – while using library flags and deterministic settings allow us to repeatedly obtain identical results across training runs on the same GPU architecture, training on different GPU architectures results in fundamentally different models.
@@ 6.3 Compression with adaptive threshold
Finally, we report the average # of $\mathsf{rev}$ corrections an $\mathrm{auditor}$ needs to perform for one training step in our two tasks (Table 3). These values are surprisingly small in comparison to the number of operations logged – only a maximum of 25, or 2e-6% (ResNet-50) and 9e-6% (GPT-2) of logged values, are actually needed by the $\mathrm{auditor}$! We also observe that severe rounding (e.g. $b=27$) completely eliminated the hardware non-determinism for our tasks, requiring no corrections from the $\mathrm{auditor}$. This shows a huge gap between the # of values currently saved by the $\mathrm{trainer}$ and those actually needed by the $\mathrm{auditor}$. An exciting direction for future work, therefore, is to develop methods that could reliably predict when a divergence will not occur and skip logging entirely, reducing the overall storage cost of our method.
@@ 6.4 Comparison with alternative approaches
7 Limitations & Future Work
Our efficient scheme for verifiable training successfully controls for hardware nondeterminism and expands the pool of potential auditors of a model training service, allowing us to envision a world where a $\mathrm{client}$ can even use two competing service providers it trusts to audit each other. However, one main limitation of our work in comparison with proof-based systems is the need for all parties to trust this third-party auditor. If the trainer is providing finetuning services on top of models that are not open-source (e.g. OpenAI), then our scheme will only work for the third-party auditors that the trainer is willing to share their model weights.
Another limitation of our work is the storage cost. While we have shown that it is significantly smaller than alternative methods, the vast majority of stored rounding decisions are not used by the auditor and are therefore unnecessary (Section 6). Therefore, an exciting direction for future work is to figure out how to mitigate this gap by better predicting when divergence between GPU computations occurs – recent work has similarly argued that a stronger profile of noise during the training process is necessary for robust verifiable training (Fang et al., 2023). Finally, other directions for future work also include adapting our method for distributed training, as well as considering adversarial manipulation of the rounding logs by the trainer.
@@ Appendix E Random Number Generation
Our verifiable training scheme requires shared randomness between the $\mathrm{trainer}$ and $\mathrm{auditor}$, which is used for deciding input data ordering and batching, weight initialization, and operations such as dropout (randomly setting outputs to zero). More formally, our scheme requires sharing the same random seed and pseudo-random generator. However, in our implementation based on pytorch (assuming the same software version between $\mathrm{trainer}$ and $\mathrm{auditor}$), we chose to rely on the the torch random seed functionality. While this successfully controls for batch input ordering and weight initialization, it is unfortunately not sufficient for random number generation, as operations such as torch.nn.randn() leverage parallelism when the requested number of values is higher than a certain amount. Specifically, we found that across T40, RTX 2080 Ti, V100, A40, and A100, given the same seed, torch.randint() produces identical tensors onlt up to size 40960. At size 40961, T40 (which is an older GPU) deviated from the rest. Likewise, at size 69633, 2080 Ti deviated from the rest, and so on. Based on these observations, we arranged for calls to torch.randint() in the dropout layer (which is the only operation using large random tensors in our tasks) to be replaced by generating and concatenating multiple random tensors of size 40960 or less. Specifically, a random tensor of size $n>40960$ is generated by concatenating ($n//40960$) random tensors of size 40960 and one random tensor of size ($n\%40960$). However, we emphasize that it is therefore important in our scheme either for both parties to implement this change a priori, or simply use an external source for pseudorandomness.

### 16 Mar 2024 v2

@@ 5.1 Accumulation Errors Start at Higher Precision Bits
Our key idea is that if nondeterminism of training between GPU types occurs due to floating-point operations, then any error will initially be introduced in the lower bits. Suppose that both $\mathrm{trainer}$ and $\mathrm{auditor}$ train at a higher floating-point (e.g. $b_{tr}=64$) precision than the client’s target model precision (e.g. $b_{m}=32$) and then periodically round (e.g. $b_{r}=32$) after intermediate computation steps (e.g. a convolution layer). One might hope that this will “erase” the errors due to nondeterminism, and prevent them from accumulating. Table 1 (Row 1) provides a toy example of how error in higher precision arithmetic is erased when rounding the result to lower precision.
@@
Unfortunately, simply rounding to the nearest FP32 after each computation during training is not sufficient for eliminating nondeterminism. The problem is due to rounding errors that straddle the rounding boundary. Consider Case A in Figure 4, which shows a divergence in the output of a computation using FP64 on two different GPUs. Because the outputs of GPU 1 and 2 are on different sides of the boundary, rounding to the nearest FP32 results in rounding to different values, introducing error. Table 1 (Row 2) provides a toy example of how error in higher precision arithmetic might transfer to the lower precision rounded result.
@@
What if the $\mathrm{trainer}$ records their rounding choice (e.g. up, down, none) for every intermediate computation? The $\mathrm{auditor}$ could then copy the $\mathrm{trainer}$’s choice, and therefore round to the exact same value and successfully control for nondeterminism. However, as we show in Case B in Figure 4, the $\mathrm{auditor}$ should not copy the $\mathrm{trainer}$’s behavior for every output. If a computation output on GPU 1 is too close to the rounded value, then it is possible that GPU 2 is also close in distance but from the opposite direction. In this case, the $\mathrm{auditor}$ should ignore the $\mathrm{trainer}$’s choice. We therefore need to introduce a threshold $\tau$ under which the $\mathrm{trainer}$ does not record their rounding choice.
@@
The success of this method relies on our ability to upper bound the divergence $d_{div}$ between any two different GPUs for any intermediate computation $f$ (i.e. difference in outputs for the same input). Let $\epsilon_{b}$ represent the distance between two FP32 values, after rounding to $b_{r}$ bits of the mantissa (Figure 3) and controlling for the exponent. We need to select $b_{r}$ and $\tau$ such that $d_{div}<\epsilon_{b_{r}}$ and $d_{div}<2\tau$ (Figure 4). Because the set of possible floating point numbers is finite, there exist optimal bounds for $b_{r}$ and $\tau$. In practice, we find that $b_{r}\leq 32$ and $\tau>0.25\cdot\epsilon_{32}$ are sufficient for standard intermediate computations in neural network training (e.g. convolution, layer norm) in FP64. We experiment with different values for $b_{r}$ in Section 6.
@@
We next discuss the specific functions that our method requires, the algorithm details for both the $\mathrm{trainer}$ and $\mathrm{auditor}$, and our approach for reducing storage costs.
@@ 5.5 Reducing storage cost
Adaptive Threshold: Recall that we need to select a threshold $\tau$ that controls for whether the $\mathrm{trainer}$ logs a rounding choice, or instead logs 1 which the $\mathrm{auditor}$ ignores. The more one can increase $\tau$, the more 1 values are recorded which, if significant, can make the rounding more compressible (due to an increase in long sequences of 1s). Furthermore, it is possible that the divergence $d_{div}$ between outputs on two different GPUs, given the same input, is function-specific. For example, while the convolution operation requires several matrix multiplications that might result in large floating point accumulation error, operations such as scaling are unlikely to result in large $d_{div}$, and a larger $\tau$ can be applied. We develop an efficient algorithm (Algorithm 4 in the Appendix) to find the optimal value for $\tau$ given a particular layer and data of output values that led to different rounding choices between any two GPUs (e.g. Case A in Figure 4). For a given rounding amount $b_{r}$ and training precision $b_{tr}$, the algorithm performs a binary search between $\tau=0.25\cdot\epsilon_{32}$ (our assumed upper bound for the $d_{div}$ between two GPUs for any function) and $\tau=0.5\cdot\epsilon_{b_{r}}$ (the rounding boundary). By performing this procedure for the different intermediate computations in a model, the $\mathrm{trainer}$ can hope to log more values of 1, leading to a stronger compression of the rounding log $F$.
@@ 6 Empirical Results
We evaluate our verifiable training method on the two large-scale models listed below with all possible $\mathrm{trainer}$ and $\mathrm{auditor}$ pairs across the following three NVIDIA GPU architectures: A40, TITAN Xp, and RTX 2080 Ti (see Appendix A for more details). Additionally, in Section 6.4, we compare our method with a proof-based system for logistic regression as no known proof-based systems exist for the large-scale models we have considered.
@@
In Figure 2 we show that nondeterminism due to GPU architecture exists for both training tasks. While using library flags and deterministic settings allow us to repeatedly obtain identical results across training runs on the same GPU architecture, training on different GPU architectures results in fundamentally different models.
@@ 6.3 Compression with adaptive threshold
Finally, we report the average number of $\mathsf{rev}$ corrections an $\mathrm{auditor}$ needs to perform for one training step in our two tasks (Table 3). These values are surprisingly small in comparison to the number of operations logged – only a maximum of 25, or 2e-6% (ResNet-50) and 9e-6% (GPT-2) of logged values, are actually needed by the $\mathrm{auditor}$! We also observe that severe rounding (e.g. $b=27$) completely eliminated the hardware non-determinism for our tasks, requiring no corrections from the $\mathrm{auditor}$. This shows a huge gap between the # of values currently saved by the $\mathrm{trainer}$ and those actually needed by the $\mathrm{auditor}$. An exciting direction for future work, therefore, is to develop methods that could reliably predict when a divergence will not occur and skip logging entirely, reducing the overall storage cost of our method.
@@ 6.4 Comparison with alternative approaches
7 Limitations and Future Work
Our scheme for verifiable training successfully controls for hardware nondeterminism and expands the pool of potential auditors of a model training service, allowing us to envision a world where a $\mathrm{client}$ can even use two competing service providers it trusts to audit each other. However, one main limitation of our work in comparison with proof-based systems is the need for all parties to trust this third-party $\mathrm{auditor}$. If the $\mathrm{trainer}$ is providing finetuning services on top of models that are not open-source (e.g. OpenAI), then our scheme will only work for the third-party auditors that the $\mathrm{trainer}$is willing to share their model weights.
Another limitation of our work is the storage cost. While we have shown that it is significantly smaller than alternative methods, the vast majority of stored rounding decisions are not used by the $\mathrm{auditor}$ and are therefore unnecessary (Section 6). Therefore, an exciting direction for future work is to figure out how to mitigate this gap by better predicting when divergence between GPU computations occurs. Recent work has similarly argued that a stronger profile of noise during the training process is necessary for robust verifiable training (Fang et al., 2023). Finally, other directions for future work also include adapting our method for distributed training, as well as considering adversarial manipulation of the rounding logs by the trainer.
@@ Appendix E Random Number Generation
Our verifiable training scheme requires shared randomness between the $\mathrm{trainer}$ and $\mathrm{auditor}$, which is used for deciding input data batching, weight initialization, and operations such as dropout (randomly setting outputs to zero). More formally, our scheme requires sharing the same random seed and pseudo-random generator. However, in our implementation based on pytorch (assuming the same software version between $\mathrm{trainer}$ and $\mathrm{auditor}$), we chose to rely on the the torch random seed functionality. While this successfully controls for batch input ordering and weight initialization, it is unfortunately not sufficient for random number generation, as operations such as torch.nn.randn() leverage parallelism when the requested number of values is higher than a certain amount. Specifically, we found that across T40, RTX 2080 Ti, V100, A40, and A100, given the same seed, torch.randint() produces identical tensors onlt up to size 40960. At size 40961, T40 (which is an older GPU) deviated from the rest. Likewise, at size 69633, 2080 Ti deviated from the rest, and so on. Based on these observations, we arranged for calls to torch.randint() in the dropout layer (which is the only operation using large random tensors in our tasks) to be replaced by generating and concatenating multiple random tensors of size 40960 or less. Specifically, a random tensor of size $n>40960$ is generated by concatenating ($n//40960$) random tensors of size 40960 and one random tensor of size ($n\%40960$). However, we emphasize that it is therefore important in our scheme either for both parties to implement this change a priori, or simply use an external source for pseudorandomness.